A 10.0g of alum is heated to drive off all of the water from the solid. Determine the mass of the dehydrated alum
1
Expert's answer
2017-04-21T12:24:05-0400
There are many alums. The solution of the problem is possible in the presence of a chemical formula. For example, let's take an alum-rich alum
M(KAl(SO₄)₂*₁₂H₂O)=39+27+32*2+16*8+12*18=474g/Mol
M(KAl(SO₄)₂)=39+27+32*2+16*8=258g/Mol
KAl(SO₄)₂*₁₂H₂O->KAl(SO₄)₂+12H₂O
474/10=258/x
X=10*258/474=5.44